OpenStudy (anonymous):

In a certain county, the number of Charter Schools is four less than twice the number of Alternative Schools. We know that there are 48 Charter Schools in the county. How many Alternative Schools are in the county

OpenStudy (moongazer):

26 alternative schools

OpenStudy (moongazer):

2x-4=48 x=52/2 x=26

OpenStudy (pfenn1):

Let x = the number of Alternative Schools.\[48=2x-4\] Solve for x.